# CurbNet: Curb Detection Framework Based on LiDAR Point Cloud Segmentation

### v1

@@ Abstract
Curb detection is an important function in intelligent driving and can be used to determine drivable areas of the road. However, curbs are difficult to detect due to the complex road environment. This paper introduces CurbNet, a novel framework for curb detection, leveraging point cloud segmentation. Addressing the dearth of comprehensive curb datasets and the absence of 3D annotations, we have developed the 3D-Curb dataset, encompassing 7,100 frames, which represents the largest and most categorically diverse collection of curb point clouds currently available. Recognizing that curbs are primarily characterized by height variations, our approach harnesses spatially-rich 3D point clouds for training. To tackle the challenges presented by the uneven distribution of curb features on the xy-plane and their reliance on z-axis high-frequency features, we introduce the multi-scale and channel attention (MSCA) module, a bespoke solution designed to optimize detection performance. Moreover, we propose an adaptive weighted loss function group, specifically formulated to counteract the imbalance in the distribution of curb point clouds relative to other categories. Our extensive experimentation on 2 major datasets has yielded results that surpass existing benchmarks set by leading curb detection and point cloud segmentation models. By integrating multi-clustering and curve fitting techniques in our post-processing stage, we have substantially reduced noise in curb detection, thereby enhancing precision to 0.8744. Notably, CurbNet has achieved an exceptional average metrics of over 0.95 at a tolerance of just 0.15m, thereby establishing a new benchmark. Furthermore, corroborative real-world experiments and dataset analyzes mutually validate each other, solidifying CurbNet’s superior detection proficiency and its robust generalizability. The code and dataset will be open sourced to: https://github.com/guoyangzhao/CurbNet/.
@@ I Introduction
Autonomous vehicles fundamentally depend on analyzing data from onboard sensors to understand their surrounding environment, a cornerstone for safe driving [1, 2]. In this context, road boundary detection is a crucial aspect of perception, delineating road and non-road areas[3]. This distinction is vital for the positioning, planning, and decision-making of self-driving cars, especially under conditions where GPS signals are obscured by trees and buildings. In such urban road environments, curbs serve as a key and effective feature for vehicle localization [4, 5]. Despite their importance, curbs, typically linear, thin, and long, pose a significant challenge for detection in complex road environments[6].
@@
Figure 1: Curb detection challenges and our proposed method. Three main challenges of curb detection are shown (a) height feature extraction (b) different density distribution of point clouds (c) Curb point cloud quantity proportion imbalance. Solution: First propose a 3D-Curb dataset. The MSCA module is designed to extract features of the xy, xz and yz planes and perform multi-feature fusion. The loss group is proposed to solve the imbalance problem. Finally, we use post-processing to further improve performance
Curb detection using LiDAR can be classified into manual feature methods and learning based methods[13]. Manual design feature methods [14, 15] typically analyze geometric relationships such as height and angle changes between adjacent points, given the difference between drivable roads and curbs [16, 17]. Most curb detection methods follow a sequential point cloud processing procedure [18, 19], including stages of candidate region extraction, manual feature setting and clustering, and post-processing for fitting estimation. These methods, due to their interpretability in terms of safety, are widely used in curb detection for autonomous driving systems. However, the reliability of these manually designed rule-based processes is limited in practical applications, as errors in early detection stages can severely impact subsequent recognition performance. Additionally, they heavily rely on manual feature design, necessitating extensive parameter adjustments for various scenarios such as straight roads, curved roads, and different types of intersections, resulting in low practical efficiency and generalization [16, 20].
As deep learning has made breakthrough applications in perception, its capabilities in automatic feature extraction and learning have significantly reduced the tedium of manual feature design and parameter adjustment[21]. Particularly, in addressing different road scenarios, it has substantially enhanced model recognition performance and robustness[22]. Recently, some researchers have explored curb detection using CNN-based methods, projecting 3D LiDAR point cloud data into 2D images from a Bird’s Eye View (BEV) perspective, followed by processing these images with CNN models to detect curbs[23, 24]. However, this direct projection of 3D LiDAR data can lead to the loss of essential spatial structural information, particularly the crucial height difference information for curb detection[25].
@@
We first propose the 3D-Curb dataset, which contains up to 7100 frames of 3D annotated point cloud data, covering a wide range of road scenes. Unlike most existing algorithms that project 3D point clouds onto 2D images, we extract features directly from 3D point clouds, preserving crucial 3D spatial information. For the challenge (a) and (b), we draw inspiration from the cylindrical voxelization approach[27], aligning the model’s input data more closely with the physical distribution of LiDAR scans. We then propose a multi-scale and multi-channel attention module, focusing not only on changes in the Z-axis height but also on enhancing the performance of multi-scale feature fusion. For the challenge (c), the difficulty of curb training in autonomous driving is objectively caused by the sparse quantity of curb point clouds. Categories related to curbs (like roads and sidewalks) can assist in the model’s detection of curbs. Addressing the severe imbalance in point cloud quantities among different categories, we introduce a new loss combination, including a adaptive cross-entropy loss and an IoU-focused loss. Finally, in order to further improve the precision of curb detection, we proposed a post-processing scheme of multi-cluster and then fitting, which effectively removes the noise around the curb detection results.
@@
    Introducing a comprehensive 3D-Curb point cloud dataset, to our knowledge which is the largest and most diverse currently available.
@@
    Proposing a new multi-scale and channel attention module and an imbalance loss strategy, based on the distribution characteristics of curb point clouds.
@@
    Implementing a multi-cluster fitting post-processing approach to further enhance detection performance.
@@ II-A Manual Feature Extraction Methods
In the early development of LiDAR technology, researchers designed mathematical functions to manually extract curb features by understanding the principles of LiDAR scanning. This process primarily involves stages of feature point extraction, feature point classification and filtering, and curb curve fitting and estimation. In [28], feature points are extracted through image segmentation and energy minimization, followed by the application of principal curves and surfaces methods in [29] for fitting detected curbs. Studies like [14, 15, 30, 31] utilize the horizontal and vertical continuity of point clouds, employing angle and height thresholds for curb feature extraction and using Gaussian Process Regression (GPR) and Random Sample Consensus for curb curve fitting. [16, 15]integrate the generalized curvature method from LOAM [23] into curb detection, refining the process with Gaussian Process Regression. [32, 16] detect curbs by analyzing ring compression in dense 3D LIDAR data, employing false positive filters and least-squares regression filters based on height values, respectively. [19, 33] propose sliding-beam segmentation and sliding-window detection methods by analyzing individual LiDAR scan lines, focusing on specific curb detection in each frame.
However, these manual feature extraction and sequential processing methods are inefficient. Not only is feature creation laborious and requires specialized knowledge, but early-stage erroneous selections can impact later detection phases, making them inadequate for complex road scenes and diverse curb shapes as noted in [34]. In this context, the advent of deep learning methods effectively solved the issues of feature extraction and generalization.
@@ II-B Deep Learning Methods
Compared to traditional methods, these approaches demonstrate robustness in various driving environments and reduce the burden of manual parameter tuning. However, converting point clouds to 2D images results in significant loss of 3D information, especially the height difference features crucial for distinguishing curbs from other categories. [38] explores voxelizing the raw 3D point cloud as a superimage input to the model, using CNNs to detect road edges and lane markings. However, 3D model recognition accuracy remains low. LCDeT [25] employs a Transformer model for curb detection in voxelized point clouds, introducing dual attention mechanisms in both temporal and spatial dimensions to ensure detection stability and accuracy. Yet, this method relies on complex model structures and high-resolution LiDAR sensors.
Our proposed method uses a single, original 3D point cloud as input, preserving more 3D feature information. Addressing the uneven distribution of LiDAR features in curb recognition scenes and the easy loss of curb height difference information, we introduce multi-scale and multi-channel attention mechanisms to enhance model recognition performance.
@@ II-C Curb Detection Datasets
There is a significant amount of research in LiDAR-based curb detection; however, high-quality datasets with curb annotations are scarce. Major automotive datasets such as NuScenes [39], KITTI [40], and SemanticKITTI [41] do not include curb annotations. The robustness of deep learning methods is closely related to the volume of data collected under various environmental conditions. These factors have somewhat hindered the application of deep learning methods in this task. [19] created a public dataset for curb detection, comprising 200 scans collected across five different scenes. [23] developed a public dataset for curb detection, including 5200 scans with Bird’s Eye View (BEV) labels, collected from urban areas. LCDeT[25] introduced a curb dataset for 128-beam LiDAR, containing 6200 frames of point clouds from urban road scenes, both during daytime and nighttime. [26] proposed a method for 3D curb detection and annotation in LiDAR point clouds, effectively reducing manual annotation time by 50%.
@@ III Methodology
Compared to other scenarios in autonomous driving, there is a notable lack of relevant curb point cloud datasets, especially those annotated for 3D point clouds. We have developed and proposed the 3D-Curb dataset based on the large-scale, open-source SemanticKITTI dataset[41], adding a new curb category while retaining the other original 28 semantic categories. This dataset was collected using a 64-line LiDAR, providing a comprehensive view of various street scenes as a universal autonomous driving dataset. In the dataset creation process, we adopted the method proposed by [34] to initially extract curb locations in a Bird’s Eye View (BEV) perspective, followed by manual annotation of representative scenes in 3D point cloud data. The 3D-Curb dataset focuses on curb annotations in the forward direction of vehicle travel, with an average range of 40.43 meters along the forward y-axis. To accentuate the curb areas, the lateral x-axis range is set to 1.3 times the road width.
The 3D-Curb dataset encompasses all scenes from sequences 00-10 in the SemanticKITTI dataset, totaling 7100 frames of finely annotated curb point cloud data. To our knowledge, this is the largest curb point cloud dataset to date and the only one annotated in 3D. Table 1 presents a comparison with other relevant curb datasets.
Figure 2: Overview of proposed CurbNet framework. From left to right, first is the input data coordinate transformation and cylindrical voxelization. Then there is the U-Net shaped encoder-decoder structure. This is followed by feature decomposition and segmentation head. Finally, the model calculates loss and outputs detection results.
@@ III-B Overview of Model Framework
As shown in Fig. 2, the CurbNet framework primarily consists of three parts: data voxelization, feature extraction, and feature decomposition and segmentation.
In the data voxelization section, to better align the input data structure to the physical characteristics of LiDAR point clouds (where the density of the circularly scanned point clouds decreases with increasing distance), we adopted a cylindrical partitioning format for point cloud voxelization, inspired by Cylinder3D[27]. We converted the Cartesian coordinates $(x,y,z)$ of the point cloud into polar coordinates (radius $\rho$ and azimuth angle $\theta$), aligning more accurately with the scanning pattern of LiDAR and preserving the point cloud’s geometric structure to the greatest extent. The voxel division was then performed uniformly across the $\rho$, $\theta$, and $z$ dimensions, ensuring larger voxels for greater distances to minimize the impact of varying point cloud densities on feature learning.
@@
In the feature extraction stage, the voxelized features of dimensions F×H×W×D were fed into an end-to-end 3D U-Net model. Diverging from the classic U-Net model’s Encoder and Decoder structure, we proposed a Multi Scale and Channel Attention Module (MSCA) tailored to the characteristic distribution of curb point clouds, employing stride=2 Sparse Convolution for pooling functions.
In the feature decomposition and segmentation stage, we decomposed the output high-dimensional features of F×H×W×D. After the training,, the feature tensor is high-rank [42], and its rich context encoding information incurs substantial modeling costs in the point cloud segmentation process. Inspired by high-rank matrix decomposition theory [43], we performed low-rank decomposition on the H×W×D dimensions using three one-dimensional convolution kernels: 3×1×1, 1×3×1, and 1×1×3, respectively. We use two layers of low-rank decomposition and add dense connections to thereby increase feature aggregation and provide more reference features for the segmentation head. The weights for each dimension were modulated through the Sigmoid function. Finally, context features were aggregated to consolidate features and output segmentation results efficiently.
III-C Multi Scale and Channel Attention (MSCA) Module
The structure of the MSCA as shown in Fig. 3, constitutes a fundamental component of both the Encoder and Decoder. Curb point clouds typically form a long curve along both sides of the road, but with the increasing scanning distance of LiDAR, the point cloud density decreases, leading to sparser curb point cloud features. This results in significant scale differences in the features represented by the same number of point clouds at different distances after voxelization. Drawing on multi-feature fusion methods[44], we uniformly blend multi-scale features derived from different convolutional kernel outputs.
The primary distinguishing features of curbs lie in the subtle height differences between the road and sidewalks, i.e., the differences along the Z-axis in the point cloud features. Unlike the Cylinder3D[27] general point cloud segmentation algorithm, which only focuses on feature learning in the xy plane and overlooks curb’s high-frequency feature information on the Z-axis, we have added a convolution module for the Z-axis D dimension. This module makes the model more attuned to curb’s high-frequency features in height differences. Inspired by text detection methods [26], we employed three asymmetrical convolution kernels to match the target areas in the xy, xz, and yz planes of the space, thereby capturing the curb feature information. Compared to the structure of two traditional 3×3×3 convolutions chained together, the MSCA module (Fig. 3) achieves thrice the efficiency in feature extraction and blending, with the same receptive field and computational resource consumption.
Figure 3: Structure of multi-scale and channel attention (MSCA) module. Three convolutions are used to match the target area on the spatial xy, xz and yz planes, and dense connections are used to obtain the multi-scale features of the curb.

### v2

@@ Abstract
Curb detection is a crucial function in intelligent driving, essential for determining drivable areas on the road. However, the complexity of road environments makes curb detection challenging. This paper introduces CurbNet, a novel framework for curb detection utilizing point cloud segmentation. To address the lack of comprehensive curb datasets with 3D annotations, we have developed the 3D-Curb dataset based on SemanticKITTI, currently the largest and most diverse collection of curb point clouds. Recognizing that the primary characteristic of curbs is height variation, our approach leverages spatially rich 3D point clouds for training. To tackle the challenges posed by the uneven distribution of curb features on the xy-plane and their dependence on high-frequency features along the z-axis, we introduce the Multi-Scale and Channel Attention (MSCA) module, a customized solution designed to optimize detection performance. Additionally, we propose an adaptive weighted loss function group specifically formulated to counteract the imbalance in the distribution of curb point clouds relative to other categories. Extensive experiments conducted on 2 major datasets demonstrate that our method surpasses existing benchmarks set by leading curb detection and point cloud segmentation models. Through the post-processing refinement of the detection results, we have significantly reduced noise in curb detection, thereby improving precision by 4.5 points. Similarly, our tolerance experiments also achieved state-of-the-art results. Furthermore, real-world experiments and dataset analyses mutually validate each other, reinforcing CurbNet’s superior detection capability and robust generalizability. The project website is available at: https://github.com/guoyangzhao/CurbNet/.
@@ I Introduction
Autonomous vehicles fundamentally depend on analyzing data from onboard sensors to understand their surrounding environment, a cornerstone for safe driving [1, 2]. In this context, road boundary detection is a crucial aspect of perception, delineating road and non-road areas[3]. This distinction is vital for the positioning, planning, and decision-making of self-driving cars, especially under conditions where GPS signals are obscured by trees and buildings. In such urban road environments, curbs serve as a key and effective feature for vehicle localization [4, 5]. Despite their importance, curbs, typically linear, thin, and long, pose significant challenges for detection in complex road environments[6].
@@
Figure 1: Curb detection challenges and our proposed method. Three main challenges of curb detection are shown (a) height feature extraction (b) different density distribution of point clouds (c) Curb point cloud quantity proportion imbalance. Solution: First propose a 3D-Curb dataset. The MSCA module is designed for multi-scale spatial feature fusion and height feature extraction. The loss group is proposed to solve the imbalance problem. Finally, we use post-processing to further improve performance.
Curb detection using LiDAR can be classified into manual feature methods and learning based methods[13]. Manual feature design methods[14, 15] typically analyze geometric relationships such as height and angle changes between adjacent points, given the difference between drivable roads and curbs [16, 17]. Most curb detection methods follow a sequential point cloud processing procedure [18, 19], including stages of candidate region extraction, manual feature setting and clustering, and post-processing for fitting estimation. These methods, due to their interpretability in terms of safety, are widely used in curb detection for autonomous driving systems. However, the reliability of these manually designed rule-based processes is limited in practical applications, as errors in early detection stages can severely impact subsequent recognition performance. Additionally, they heavily rely on manual feature design, necessitating extensive parameter adjustments for various scenarios such as straight roads, curved roads, and different types of intersections, resulting in low practical efficiency and generalization [16, 20].
With the breakthroughs in deep learning applications in perception, its capabilities in automatic feature extraction and learning have significantly reduced the tedium of manual feature design and parameter adjustment[21]. Particularly, in addressing different road scenarios, it has substantially enhanced model recognition performance and robustness[22]. Recently, some researchers have explored curb detection using CNN-based methods, projecting 3D LiDAR point cloud data into 2D images from a Bird’s Eye View (BEV) perspective, followed by processing these images with CNN models to detect curbs[23, 24]. However, this direct projection of 3D LiDAR data can lead to the loss of essential spatial structural information, particularly the crucial height difference information for curb detection[25].
@@
We first propose the 3D-Curb dataset, which contains 3D annotated point cloud data with a wide range of road scenes. Unlike most existing algorithms that project 3D point clouds onto 2D images, we extract features directly from point clouds, preserving crucial 3D spatial information. To address challenges (a) and (b), we propose a Multi-Scale and Channel Attention (MSCA) module. This module employs a multi-scale fusion stage to mitigate the uneven distribution of curb point clouds and a channel attention stage to dynamically capture height variations along the z-axis feature. For challenge (c), which pertains to the severe imbalance in point cloud quantities among different categories, we introduce a novel loss combination. This includes an adaptive cross-entropy (ACE) loss and an IoU-focused loss, designed to handle the imbalanced data distribution effectively. Moreover, recognizing the inherent sparsity of curb point clouds, we augment the training process by adding curb-related categories (such as roads and sidewalks) to assist in model learning. Finally, to further improve the precision of curb detection, we propose a post-processing scheme of multi-cluster and then fitting, which effectively removes the noise around the curb detection results.
@@
    Introducing a comprehensive 3D-Curb point cloud dataset based on SemanticKITTI, to our knowledge which is the largest and most diverse currently available.
@@
    Proposing a new Multi-Scale and Channel Attention (MSCA) module and an imbalance loss strategy, tailored to the distribution characteristics of curb point clouds.
@@
    Developing a multi-cluster fitting post-processing approach to further enhance detection performance.
@@ II-A Manual Feature Extraction Methods
In the early development of LiDAR technology, researchers designed mathematical functions to manually extract curb features by understanding the principles of LiDAR scanning. This process primarily involves stages of feature point extraction, feature point classification and filtering, and curb curve fitting and estimation. In [27], feature points are extracted through image segmentation and energy minimization, followed by the application of principal curves and surfaces methods in [28] for fitting detected curbs. Studies like [14, 15, 29, 30] utilize the horizontal and vertical continuity of point clouds, employing angle and height thresholds for curb feature extraction and using Gaussian Process Regression (GPR) and Random Sample Consensus for curb curve fitting. [16, 15]integrate the generalized curvature method from LOAM [23] into curb detection, refining the process with GPR. [31, 16] detect curbs by analyzing ring compression in dense 3D LIDAR data, employing false positive filters and least-squares regression filters based on height values, respectively. [19, 32] propose sliding-beam segmentation and sliding-window detection methods by analyzing individual LiDAR scan lines, focusing on specific curb detection in each frame.
However, these manual feature extraction and sequential processing methods are inefficient. Not only is feature creation laborious and requires specialized knowledge, but early-stage erroneous selections can impact later detection phases, making them inadequate for complex road scenes and diverse curb shapes[33]. In this context, deep learning methods can effectively solve the issues of feature extraction and generalization.
@@ II-B Deep Learning Methods
Compared to traditional methods, these approaches demonstrate robustness in various driving environments and reduce the burden of manual parameter tuning. However, converting point clouds to 2D images results in significant loss of 3D information, especially the height difference features crucial for distinguishing curbs from other categories. [37] explores voxelizing the raw 3D point cloud as a superimage input to the model, using CNNs to detect road edges and lane markings, but its recognition accuracy remains low. LCDeT [25] employs a Transformer model for curb detection in voxelized point clouds, introducing dual attention mechanisms in both temporal and spatial dimensions to ensure detection stability and accuracy. Yet, this method relies on complex model structures and high-resolution LiDAR sensors.
Our proposed method uses original point cloud as input, preserving more 3D feature information. Addressing the uneven distribution of curb point cloud feature and the easy loss of curb height difference information, we introduce a multi-scale and channel attention mechanisms to enhance performance.
@@ II-C Curb Detection Datasets
There is a significant amount of research in LiDAR-based curb detection; however, high-quality datasets with 3D annotations are scarce. Major automotive datasets such as NuScenes [38], KITTI [39], and SemanticKITTI [40] do not include curb annotations. The robustness of deep learning methods is closely related to the volume of data collected under various environmental conditions. These factors have somewhat hindered the application of deep learning methods in this task. [19] created a public dataset for curb detection, comprising 200 scans collected across five different scenes. [23] developed a public dataset for curb detection, including 5200 scans with BEV labels, collected from urban areas. LCDeT[25] introduced a curb dataset for 128-beam LiDAR, containing 6200 frames of point clouds from urban road scenes, both during daytime and nighttime. [26] proposed a method for 3D curb detection and annotation in LiDAR point clouds, effectively reducing manual annotation time by 50%.
@@
Figure 2: 3D-Curb dataset construction process. Mainly developed based on the standard SemanticKITTI dataset.
@@ III Methodology
Compared to other autonomous driving scenarios, there is a significant lack of relevant curb point cloud datasets, especially those with 3D annotations. Building on the large-scale open-source SemanticKITTI dataset [40], we have developed and introduced the 3D-Curb dataset. This dataset retains the original 28 semantic categories while adding a new curb category. It was collected using a Velodyne HDL-64E LiDAR, providing comprehensive views of various street scenes as a general-purpose autonomous driving dataset.
The construction process of our dataset is illustrated in Fig. 2. Due to the high-frequency data acquisition of the SemanticKITTI dataset, the high similarity between frames can lead to model overfitting and significantly increase the annotation workload. Therefore, we randomly selected 7100 representative frames from sequences 00-10. We utilized the high-quality road annotations provided by the SemanticKITTI dataset and applied the Ground Plane Fitting method proposed by [41] to extract the boundaries within the road labels. We then extended these parameters to obtain the curb labels. However, due to static/dynamic obstacles and other occlusions, these automatically generated curb annotations contained many inaccuracies. Thus, we manually refined the curb labels in the BEV perspective to ensure high-quality annotations. The 3D-Curb dataset focuses on curb annotations in the forward direction of vehicle travel, with an average range of 40.43 meters along the forward y-axis. To accentuate the curb areas, the lateral x-axis range is set to 1.3 times the road width.
To the best of our knowledge, this is the largest curb point cloud dataset to date and the only one with 3D annotations. Table. I compares our dataset with other related curb datasets.
Figure 3: Overview of proposed CurbNet framework. From left to right, first is point cloud data input and voxelization. Then there is a 5-layer deep encoder-decoder structure. Next comes the feature aggregation and segmentation head. Finally, the post-processing refinement of the detection results.
@@ III-B Overview of Model Framework
As illustrated in Fig. 3, the CurbNet framework comprises four main components: point cloud input with voxelization, feature extraction, feature aggregation to segmentation, and post-processing refinement.
In most road scenarios, curbs are primarily located at the junction between the road and the sidewalk. To enhance the model’s ability to learn curb features, we introduced additional labels for the road and sidewalk during the training process. To align the input point cloud data with physical space characteristics (where the density of circularly scanned point clouds decreases with increasing distance), we partition the voxel blocks based on the point cloud distance. This approach minimizes the impact of uneven point cloud distribution.
@@
During the feature extraction stage, voxelized features with dimensions $F\times H\times W\times D$ are fed into a 5-layer deep encoder-decoder structure. Each encoder and decoder backbone incorporates the Multi-Scale and Channel Attention (MSCA) module, specifically designed for curb point cloud feature extraction. A stride-2 sparse convolution is employed as the pooling function.
Subsequently, the high-dimensional voxelized features ($F\times H\times W\times D$) are aggregated and then input into the segmentation head to obtain the detection results. Finally, the curb detection results undergo post-processing involving multi-clustering, fitting, and noise removal to further enhance detection accuracy.
The integration of these components within the CurbNet framework enables robust and precise curb detection, addressing the challenges posed by varying point cloud densities and the complex spatial characteristics of road environments.
III-C Multi-Scale and Channel Attention (MSCA) Module
The structure of the MSCA module, illustrated in Fig. 4, serves as a fundamental component of the encoder-decoder architecture. The MSCA module is composed of two main parts: multi-scale fusion and channel attention.
Multi-Scale Fusion. Curbs in point clouds typically form a long curve along both sides of the road. However, as the LiDAR scanning distance increases, the point cloud density decreases, leading to sparser curb features. This results in significant scale differences in the features represented by the same number of point clouds at different distances after voxelization. To address this, we designed the MSCA module with a multi-scale fusion strategy. We employ convolution groups with different strides (1, 3, 5) to construct a feature pyramid, and then use dense pyramid connections to deeply fuse the output features of different scales. Inspired by text detection methods [42], we adopt three asymmetric sparse convolution kernels to target the regions in the xy, xz, and yz planes within the voxel space. This approach captures multi-dimensional curb feature information. Compared to traditional 3×3×3 convolutions, the combination of asymmetric sparse convolutions offers higher computational efficiency while maintaining the same receptive field.
Channel Attention. The primary distinguishing feature of curbs is the subtle height difference between the road and the sidewalk, reflected in the z-axis of point cloud data. Unlike general point cloud segmentation algorithms such as Cylinder3D [43], which primarily focus on feature learning in the xy plane, we designed the channel attention module to capture high-frequency features along the z-axis. Initially, a 1×1×D sparse convolution is applied to preliminarily extract channel features. This is followed by an MLP to encode and decode the channel features, enhancing the representation of height features. The output of the MLP is then passed through a softmax function to generate dynamic weights for each channel. These dynamic weights are multiplied with the channel features to produce the final channel attention output.
By combining multi-scale feature extraction with dynamic channel attention, the MSCA module effectively captures both spatial and height variations in point cloud data, thereby enhancing the model’s ability to detect curbs.
Figure 4: Structure of multi-scale and channel attention (MSCA) module. S Conv means Sparse convolution layer. Multi-Scale Fusion is mainly used to fuse spatial features of different scales, and Channel Attention is used to dynamically extract height features of the z-axis.